### v1

# Multi-view Hypergraph-based Contrastive Learning Model for Cold-Start Micro-video Recommendation thanks: * indicates corresponding author

Abstract
With the widespread use of mobile devices and the rapid growth of micro-video platforms such as TikTok and Kwai, the demand for personalized micro-video recommendation systems has significantly increased. Micro-videos typically contain diverse information, such as textual metadata, visual cues (e.g., cover images), and dynamic video content, significantly affecting user interaction and engagement patterns. However, most existing approaches often suffer from the problem of over-smoothing, which limits their ability to capture comprehensive interaction information effectively. Additionally, cold-start scenarios present ongoing challenges due to sparse interaction data and the underutilization of available interaction signals.
To address these issues, we propose a Multi-view Hypergraph-based Contrastive learning model for cold-start micro-video Recommendation (MHCR). MHCR introduces a multi-view multimodal feature extraction layer to capture interaction signals from various perspectives and incorporates multi-view self-supervised learning tasks to provide additional supervisory signals. Through extensive experiments on two real-world datasets, we show that MHCR significantly outperforms existing video recommendation models and effectively mitigates cold-start challenges. Our code is available at https://anonymous.4open.science/r/MHCR-02EF.
Index Terms:
Micro-video recommendation, Cold-start problem, Multimodal feature extraction, Hypergraph model, Self-supervised learning
I Introduction
In recent years, the rapid expansion of platforms such as TikTok and Kwai has heightened the need for effective personalized micro-video recommendation systems. Unlike other content types, such as news or music, micro-videos involve richer multi-modal features, typically including titles (text), cover images (visual), and the videos themselves (visual). These multi-modal features are crucial in shaping users’ behavioral decisions and largely determine whether a user engages with a micro-video [1, 2, 3].
As a specialized area of recommendation systems, video recommendation has made numerous valuable advancements. For instance, MMGCN [4] integrates multi-modal features into a user–item bipartite graph neural network to refine user–item interactions, mitigating the impact of spurious interactions on model performance. MMGCL [1] introduces a novel negative sampling technique to learn cross-modal correlations, ensuring each modality contributes effectively and enhancing recommendation accuracy. CMI [2] leverages contrastive learning across multiple interests to capture users’ diverse preferences, improving the robustness of interest embeddings.
Despite the progress made by existing micro-video recommendation methods, they still face significant challenges in cold-start scenarios, primarily due to two issues:
  • Sparse interaction signals: Micro-video recommendation is heavily impacted by the long-tail distribution, where most micro-videos receive minimal interaction signals. This results in suboptimal performance of current models in cold-start conditions.
  • Underutilization of interaction information: Most existing methods rely on Graph Neural Networks (GNNs) to aggregate interaction information through message passing. However, these methods often suffer from over-smoothing, limiting the model’s ability to capture comprehensive interaction information.
To address these challenges, we propose a Multi-view Hypergraph-based Contrastive Learning Model for Cold-Start Micro-video Recommendation (MHCR). First, MHCR employs hypergraphs to capture more extensive interaction information, facilitating the effective propagation of interaction signals. Second, MHCR incorporates multiple self-supervised learning tasks to provide additional supervision signals, enhancing the model’s learning capability. Our key contributions are as follows:
  • We introduce MHCR, the first model to leverage hypergraphs and contrastive learning to tackle the cold-start problem in micro-video recommendation.
  • We design a multi-view multimodal feature extraction layer to repeatedly harness interaction information from multiple perspectives and introduce multi-view self-supervised learning tasks to provide auxiliary supervision signals.
  • Extensive experiments on two real-world datasets validate the effectiveness of MHCR and ablation studies confirm the contribution of each module within the model.
Figure 1: Overview of the MHCR Architecture.
II PROPOSED METHOD
II-A Problem Definition
Let $U=\{u\}$ and $I=\{i\}$ represent the sets of users and items, respectively. The corresponding ID embeddings are represented as $E_{ui}\in\mathbb{R}^{d\times(|U|+|I|)}$, where $d$ denotes the embedding dimension. Additionally, we define $M=\{i,v,t\}$, which encompasses the image, video, and text modalities. For each modality $m\in M$, the item modality features are represented as $E_{i,m}\in\mathbb{R}^{d_{m}\times|I|}$. These features are projected into a unified vector space by pre-trained models (e.g., ViT for images, V-Swin for videos, and MiniLM for text), and multimodal features are extracted from multi-view. The framework is shown in Fig. 1.
II-B Multi-view Multimodal Feature Extraction Layer
User–Item Graph Feature Extraction Layer. Inspired by LightGCN, we design a user–item graph aimed at effectively capturing high-order collaborative signals, and the representation at the $l$-th layer is defined as follows:
$E_{ui}^{(l)}=\sum_{i\in N_{u}}\frac{1}{\sqrt{|N_{u}|\cdot|N_{i}|}}E_{ui}^{(l-1)},$ (1)
where $N_{u}$ denotes the neighbors of user $u$, and $N_{i}$ denotes the neighbors of item $i$.
The final user–item representation is obtained by summing the embeddings from all layers: $E_{ui}=\sum_{l=0}^{L}E_{ui}^{(l)}$, where $L$ is the total number of graph convolution layers.
Item–Item Graph Feature Extraction Layer. Similar to the user–item view, we construct an item–item affinity graph to capture modality-specific correlations, and the affinity between items $a$ and $b$ in modality $m$ is defined as follows:
$s_{a,b}^{m}=\frac{(e_{a}^{m})^{\text{T}}\cdot e_{b}^{m}}{\|e_{a}^{m}\|\|e_{b}^{m}\|},$ (2)
where $e_{a}^{m}$ and $e_{b}^{m}$ are the modality features of items $a$ and $b$.
We apply KNN sparsification to keep the top $K$ nearest neighbors for each item and normalize the affinity matrix. The item modality features ($E_{i,m}$) are then propagated through the graph. The overall item embedding is computed as: $E_{ii}=\sum_{m}S_{m}E_{i,m}$.
TABLE I: Performance Metrics of Different Models on Two MicroLens Datasets
Dataset Metric YouTube VBPR MMGCN LightGCN GRCN LayerGCN BM3 Freedom MGCN MHCR
MicroLens-50K R@10 0.0375 0.0544 0.0403 0.0365 0.0631 0.0627 0.0565 0.0656 0.0708 0.0736
R@20 0.0632 0.0888 0.067 0.0534 0.0982 0.0994 0.0918 0.1028 0.1089 0.1102
N@10 0.0178 0.0273 0.0197 0.0284 0.0328 0.032 0.0281 0.0334 0.0363 0.0383
N@20 0.0245 0.0361 0.0264 0.0345 0.0415 0.0414 0.0372 0.0429 0.0459 0.0477
MicroLens-100K R@10 0.0392 0.0624 0.0405 0.0388 0.0682 0.0730 0.0601 0.0654 0.0717 0.0798
R@20 0.0648 0.1002 0.0678 0.056 0.1057 0.1120 0.0975 0.1016 0.1096 0.1187
N@10 0.0188 0.0314 0.0202 0.0306 0.0353 0.0382 0.0305 0.0337 0.0371 0.042
N@20 0.0252 0.0410 0.0271 0.0367 0.0448 0.0480 0.0401 0.0431 0.0467 0.0519
Hypergraph High-Level Feature Extraction Layer. To capture higher-order dependencies between users, items, and their attributes, we introduce learnable hyperedge embeddings: $H_{i}^{m}=E_{i}^{m}\cdot V_{m}^{\top}$, $H_{u}^{m}=X_{u}\cdot(H_{i}^{m})^{\top}$, where $E_{i}^{m}$ is the item feature matrix for modality $m$, $V_{m}$ is the hyperedge vector matrix, and $X_{u}$ is derived from the user-item interaction matrix $X$. Hypergraph message passing facilitates global information transfer by using hyperedges:
$\begin{cases}E_{i}^{m,h+1}=DROP(H_{i}^{m})\cdot DROP((H_{i}^{m})^{\top})\cdot E_{i}^{m,h},\\ E_{u}^{m,h+1}=DROP(H_{u}^{m})\cdot DROP((H_{i}^{m})^{\top})\cdot E_{i}^{m,h},\end{cases}$ (3)
The hypergraph embedding matrix $E_{h}$ is obtained by aggregating the embeddings across all modalities:
$E_{h}=\sum_{m}E_{m}^{H},\quad E_{m}^{H}=[E_{u}^{m},E_{i}^{m}].$
By concatenating the user–item features $E_{ui}$, item–item features $E_{ii}$, and hypergraph embedding $E_{h}$, the final modality feature $E^{*}\in\mathbb{R}^{d\times(|U|+|I|)}$ is obtained.
II-C Multi-view Self-Supervised Learning
To ensure the effective fusion of global embeddings across various modalities, we introduce a cross-modal hypergraph contrastive learning:
$L_{hc}=\sum_{x\in U\cup I}-\log\frac{\sum_{m}\exp\left(\frac{s(E_{x}^{m,H},E_{x}^{m^{\prime},H})}{\tau}\right)}{\sum_{x^{\prime}\in U\cup I}\sum_{m}\exp\left(\frac{s(E_{x}^{m,H},E_{x}^{t,H})}{\tau}\right)},$ (4)
where $s(\cdot)$ denotes the cosine similarity; $\exp$ denotes the exponential function; and $\tau$ denotes the temperature parameter.
To align graph and hypergraph embeddings, a graph-hypergraph contrastive learning strategy is used, and the loss is defined as follows:
$L_{ghc}=\sum\left(-\log\frac{\exp\left(\frac{s(E^{G},E^{H})}{\tau}\right)}{\exp\left(\frac{s(E^{G},E^{H})}{\tau}\right)+\sum\exp\left(\frac{s(E^{G},E^{H})}{\tau}\right)}\right).$ (5)
Finally, the overall loss combines the BPR loss $L_{BPR}$ with the hypergraph contrastive loss $\lambda_{hc}L_{hc}$ and the graph hypergraph contrastive loss $\lambda_{ghc}L_{ghc}$, ensuring a balance between ranking and contrastive learning.
II-D Prediction
Based on the refined behavioral and multimodal features, the final representations for users and items are formulated as follows: $e_{u}=e_{u,ui}+e_{u,ii}+e_{u,h},\quad e_{i}=e_{i,ui}+e_{i,ii}+e_{i,h}$. The interaction likelihood between user $u$ and item $i$ is determined by computing their inner product:
$f_{predict}(u,i)=\hat{y}_{ui}=e_{u}^{\top}e_{i}.$ (6)
III EXPERIMENTS
In this section, we conduct experiments to assess the effectiveness of the proposed model across two MicroLens datasets. The experimental results provide clear insights into the following research questions: (1) RQ1: How does the MHCR’s performance compare to existing video recommendation techniques? (2) RQ2: What is the impact of the individual components on the overall performance of the MHCR? (3) RQ3: How do variations in hyper-parameter configurations affect the MHCR’s outcomes? (4) RQ4: How does the performance of the MHCR model in cold-start scenarios compare to baseline models?
III-A Experimental Settings
III-A1 Datasets
In this study, we utilized two datasets from the MicroLens series, namely MicroLens-50K and MicroLens-100K. The MicroLens-50K dataset comprises 50,000 users, 19,220 items, and a total of 359,708 interaction records, resulting in an average of 7.19 interactions per user and 18.71 interactions per item, with an overall sparsity level of 99.96%. Similarly, the MicroLens-100K dataset includes 100,000 users, 19,738 items, and 719,405 interaction records. Beyond the interaction data, the MicroLens datasets also feature 15,580 tags representing detailed video categories. The length of interaction sequences typically falls between 5 and 15, with most micro-videos lasting less than 400 seconds, providing a more comprehensive contextual dataset for in-depth analysis.
III-A2 Compared methods
To verify the effectiveness of the proposed MHCR, we conducted a detailed comparison of three different types of recommendation models: (1) General Deep Learning Recommendation Models, including YouTube[5] and VBPR[6]; (2) Graph-based Recommendation Models, including LightGCN[7] and LayerGCN[8]; and (3) Multimodal Graph-based Recommendation Models, including MMGCN[4], GRCN[9], BM3[10], Freedom[11], and MGCN[12].
III-A3 Evaluation Protocols
To evaluate model performance, we randomly split each user’s interaction history into 70% for training, 10% for validation, and 20% for testing. Recall@$n$ and NDCG@$n$ (with $n\in\{10,20\}$) are used as evaluation metrics to assess the effectiveness of the top-$n$ recommendations. In the following discussions, we use R@$n$ and N@$n$ as shorthand for Recall@$n$ and NDCG@$n$, respectively.
III-B Overall Performance (RQ1)
Based on the results presented in Table I, we compare the performance of the proposed MHCR model with baseline models and highlight the following key findings:
  1. 1.
    The MHCR model consistently outperforms all baseline models across the MicroLens-50K and MicroLens-100K datasets. Specifically, on the MicroLens-50K dataset, MHCR demonstrates a significant improvement of 3.96% in Recall@10 and 5.51% in NDCG@10 when compared to the closest baseline model, MGCN. Similarly, on the larger MicroLens-100K dataset, MHCR outperforms other models with an increase of 11.30% in Recall@10 and 13.19% in NDCG@10. These results highlight the effectiveness of the proposed Hypergraph Embedding Module and the Self-Supervised Learning mechanism, which allow MHCR to capture better the sparse interaction patterns inherent in micro-video recommendation tasks.
  2. 2.
    Multimodal Graph-based Recommendation Models generally outperform Graph-based Recommendation Models, underscoring the importance of incorporating multimodal information, which plays a pivotal role in enhancing the predictive performance of recommendation systems.
  3. 3.
    Graph-based Recommendation Models consistently surpass general deep learning recommendation models, highlighting the capability of Graph Neural Networks to effectively aggregate neighborhood information, thus enabling more precise modeling of user and item representations.
III-C Ablation Study(RQ2)
To rigorously evaluate the contribution of each component in our proposed MHCR model, we conducted an ablation study by systematically disabling each module and analyzing the impact on performance. The experiments were conducted on the MicroLens-100K dataset using the following configurations: (1) w/o UI: removing the User–Item Graph Embedding Module. (2) w/o II: removing the Item–Item Graph Embedding Module. (3) w/o HEM: removing the Hypergraph Embedding Module. (4) w/o HC: removing the Hypergraph Contrastive Module. (5) w/o GHC: removing the Graph-Hypergraph Contrastive Module.
Fig. 2 illustrates the results of our ablation study on the MicroLens-100K. The original MHCR configuration consistently outperforms its modified versions, underscoring the critical importance of each component. The performance drop observed when individual modules are removed respectively highlights the key roles of the graph, hypergraph, and self-supervised learning components, confirming their important contributions to the overall effectiveness of the MHCR model.
Figure 2: Ablation Study Results of MHCR.
III-D Sensitivity Analysis (RQ3)
In this section, we perform a series of experiments to identify the optimal hyperparameters, explicitly focusing on the number of hypergraph layers ($\boldsymbol{hyper\_num}$), the weight of the hypergraph contrastive loss ($\boldsymbol{\lambda_{hc}}$), and the weight of the graph-hypergraph contrastive loss ($\boldsymbol{\lambda_{ghc}}$).
The results presented in Fig.3 reveal three key insights: First, setting $hyper\_num$ to 32 yields the best performance while increasing the number of layers slightly degrades performance due to increased complexity. Second, the optimal value for $\lambda_{hc}$ is found to be $1.00\times 10^{-5}$, which enhances the effectiveness of contrastive learning without leading to overfitting. Finally, $\lambda_{ghc}$ performs optimally at 0.01, effectively aligning representations, while higher values result in diminished performance.
Figure 3: Effect of Hypergraph Parameters and Contrastive Learning on Recommendation Performance.
III-E Results on Cold-start Scenarios (RQ4)
In the cold-start scenario, where users have fewer than three interactions, the experimental results across the two datasets demonstrate that the MHCR model consistently outperforms the baselines, as shown in Table II. On the MicroLens-50K dataset, MHCR improves Recall@10 and NDCG@10 by 4.8% and 7.3%, respectively. Similarly, on the MicroLens-100K dataset, MHCR improves by 5.0% in Recall@20 and 6.8% in NDCG@20. These results highlight MHCR’s effectiveness in addressing cold-start challenges through its multi-view hypergraph contrastive learning approach.
TABLE II: Performance Comparison of Different Models for Cold-Start Users on Two MicroLens Datasets
Dataset Metric GRCN LayerGCN BM3 Freedom MGCN MHCR
M-50K R@10 0.0521 0.0527 0.0471 0.053 0.0588 0.0616
R@20 0.081 0.0826 0.0789 0.0839 0.0912 0.0937
N@10 0.0267 0.0267 0.0237 0.0267 0.0299 0.0321
N@20 0.034 0.0341 0.0316 0.0344 0.038 0.0396
M-100K R@10 0.0562 0.0588 0.0485 0.0544 0.0605 0.0655
R@20 0.0874 0.092 0.0807 0.0882 0.0937 0.0984
N@10 0.0288 0.0299 0.024 0.0275 0.0311 0.0342
N@20 0.0366 0.0384 0.0321 0.0359 0.0396 0.0423
IV Conclusion
To address the cold-start issue and limitations in capturing comprehensive interaction information, this paper presents MHCR, which leverages hypergraph structures for intricate interaction patterns and self-supervised learning to improve predictive performance. It enhances embeddings through user–item and item–item graphs for multi-hop connections and intra-modal relationships. Hypergraph embeddings capture higher-order associations, while contrastive losses boost robustness and distinctiveness across modalities. Experiments on real-world datasets confirm MHCR’s effectiveness in handling multimodal data and cold-start challenges, paving the way for further improvements in hypergraph-based recommendation systems.

### v2

@@
-Multi-view Hypergraph-based Contrastive Learning Model for Cold-Start Micro-video Recommendation thanks: * indicates corresponding author
+Multi-view Hypergraph-based Contrastive Learning Model for Cold-Start Micro-video Recommendation
 Abstract
@@
 Through extensive experiments on two real-world datasets, we show that MHCR significantly outperforms existing video recommendation models and effectively mitigates cold-start challenges.
-Our code is available at https://anonymous.4open.science/r/MHCR-02EF.
+Our code is available at https://github.com/sisuolv/MHCR.
 Index Terms:
@@
 Unlike other content types, such as news or music, micro-videos involve richer multi-modal features, typically including titles (text), cover images (visual), and the videos themselves (visual).
-These multi-modal features are crucial in shaping users’ behavioral decisions and largely determine whether a user engages with a micro-video [1, 2, 3].
+These multi-modal features are crucial in shaping users’ behavioral decisions and largely determine whether a user engages with a micro-video [1, 2, 3, 4, 5].
 As a specialized area of recommendation systems, video recommendation has made numerous valuable advancements.
@@
 $E_{ui}^{(l)}=\sum_{i\in N_{u}}\frac{1}{\sqrt{|N_{u}|\cdot|N_{i}|}}E_{ui}^{(l-1)},$ (1)
-where $N_{u}$ denotes the neighbors of user $u$, and $N_{i}$ denotes the neighbors of item $i$.
+where $N_{u}$ and $N_{i}$ denote the neighbors of user $u$ and item $i$, respectively.
 The final user–item representation is obtained by summing the embeddings from all layers: $E_{ui}=\sum_{l=0}^{L}E_{ui}^{(l)}$, where $L$ is the total number of graph convolution layers.
@@
 Hypergraph High-Level Feature Extraction Layer.
-To capture higher-order dependencies between users, items, and their attributes, we introduce learnable hyperedge embeddings: $H_{i}^{m}=E_{i}^{m}\cdot V_{m}^{\top}$, $H_{u}^{m}=X_{u}\cdot(H_{i}^{m})^{\top}$, where $E_{i}^{m}$ is the item feature matrix for modality $m$, $V_{m}$ is the hyperedge vector matrix, and $X_{u}$ is derived from the user-item interaction matrix $X$.
+To capture higher-order dependencies between users, items, and their attributes, we introduce learnable hyperedge embeddings: $H_{i}^{m}=E_{i}^{m}\cdot V_{m}^{\top}$, $H_{u}^{m}=X_{u}\cdot(H_{i}^{m})^{\top}$, where $E_{i}^{m}$ is the item feature matrix for modality $m$, and $V_{m}\in\mathbb{R}^{K\times d_{m}}$ represents the hyperedge vector matrix, with $K$ denoting the number of hyperedges.
+Additionally, $X_{u}$ is derived from the user-item interaction matrix $X$.
 Hypergraph message passing facilitates global information transfer by using hyperedges:
@@
 III-D Sensitivity Analysis (RQ3)
-In this section, we perform a series of experiments to identify the optimal hyperparameters, explicitly focusing on the number of hypergraph layers ($\boldsymbol{hyper\_num}$), the weight of the hypergraph contrastive loss ($\boldsymbol{\lambda_{hc}}$), and the weight of the graph-hypergraph contrastive loss ($\boldsymbol{\lambda_{ghc}}$).
+In this section, we perform a series of experiments to identify the optimal hyperparameters, explicitly focusing on the number of hypergraph layers ($\boldsymbol{hyper\_num}$), the ratio of the hypergraph contrastive loss ($\boldsymbol{\lambda_{hc}}$), and the ratio of the graph-hypergraph contrastive loss ($\boldsymbol{\lambda_{ghc}}$).
 The results presented in Fig.3 reveal three key insights: First, setting $hyper\_num$ to 32 yields the best performance while increasing the number of layers slightly degrades performance due to increased complexity.